OpenStudy (anonymous):

Use basic identities to simplify the expression. cos θ - cos θ sin^2θ

OpenStudy (anonymous):

okay, I assume you mean this :\[\cos \theta - \cos \theta \sin^{2}\theta\]

OpenStudy (anonymous):

yes

OpenStudy (cggurumanjunath):